Check my work Check My Work button is now enabledItem 3 Item 3 35 points Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost. Last year, the company sold 60,000 of these balls, with the following results: Sales (60,000 balls) $ 1,500,000 Variable expenses 900,000 Contribution margin 600,000 Fixed expenses 375,000 Net operating income $ 225,000 Required: 1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level. 2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls? 3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year? 4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs? 5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls? 6. Refer to the data in (5) above. a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year? b. Assume the new plant is built and that next year the company manufactures and sells 60,000 balls (the same number as sold last year). Prepare a contribution format income statement and Compute the degree of operating leverage.
Answer #1

1. Compute

(a) last year's CM ratio and the break-even point in balls

CM per unit= Sales price per unit-Variable expenses per unit

= $25-15= $10

CM ratio= CM per unit*100/Sales price per unit

= $10*100/25= 40%

Break even point in balls= Fixed cost/CM per unit

= $375000/10= 37500 balls

(b) the degree of operating leverage at last year’s sales level.

Degree of operating leverage= Contribution margin/Net operating income

= $600000/225000= 2.67

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

New variable expenses per ball= $15+3= $18

New contribution margin per ball= $25-18= $7

CM ratio= Contribution margin*100/Sales

= $7*100/25= 28%

Break even point in balls= Fixed cost/Contribution margin per ball

= $375000/7= 53571 balls

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year?

Sales in balls= (Fixed cost+Target income)/Contribution margin per ball

= ($375000+225000)/7= 85714 balls

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

Selling price= Variable expense+Contribution margin

Let selling price= X

Then, contribution margin= 40%*x= 0.4x

x= $18+0.4x

0.6x= $18

x= $30

Selling price= x= $30 per ball

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

New variable expenses= $15*60%= $9

Contribution margin per ball= Selling price-Variable expense

= $25-9= $16

Contribution margin ratio= Contribution margin*100/Sales

= $16*100/25= 64%

New fixed expenses= $375000*2= $750000

Break even point in balls= Fixed cost/Contribution margin per unit

= $750000/16= 46875 balls

6. Refer to the data in (5) above. a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $225,000, as last year?

Break eve point in balls= (Fixed cost+Target income)/Contribution margin per unit

= ($750000+225000)/16= 60938 balls

b. Assume the new plant is built and that next year the company manufactures and sells 60,000 balls (the same number as sold last year). Prepare a contribution format income statement and Compute the degree of operating leverage.

Northwood Company
Contribution Income Statement
Sales (60000*$25) $1500000
Variable expenses (60000*$9) -540000
Contribution margin 960000
Fixed expenses -750000
Net operating income $210000
Degree of operating leverage (960000/210000) 4.57